# Adaptive variational quantum dynamics simulations with compressed circuits and fewer measurements

### v1

@@ III Models
In addition to AVQDS and AVQDS(T), we also simulated quantum dynamics with the first-order Trotter decomposition Trotter (1959); Nielsen and Chuang (2011) and VQDS with a fixed ansatz, namely the Hamiltonian variational ansatz (HVA) Wecker et al. (2015b). In Trotter decomposition, the time evolution of a quantum state over a single time step $\delta t$ is approximated as |$\ket{\Psi(t+\delta t)}\approx\prod\limits_{\mu}e^{-i\delta tj_{\mu}\,\hat{h}_{\mu}}\ket{\Psi(t)}$, where $j_{\mu}\hat{h}_{\mu}$ are the terms in the Hamiltonian: ^$\hat{\mathcal{H}}=\sum\limits_{\mu}j_{\mu}\hat{h}_{\mu}$. $\hat{h}_{\mu}$ are usually Pauli strings and $j_{\mu}$ are constants. A basic feature of such Trotter circuits is that the circuit depth keeps increasing with the simulation time $t$ as the circuit depth scales linearly with $t$. Motivated by the form of these Trotter circuits, the HVA entails parameterized unitary gates generated by the individual Pauli terms $\hat{h}_{\mu}$. The ansatz uses a fixed number of layers, $L$, and the rotation angles entering the parameterized unitaries are allowed to vary in space and from layer to layer. Thus, the variational quantum state within HVA can be written as
@@ IV Results and Discussions
IV.1 Technical approaches for solving the equations of motion
An essential process in both VQDS and AVQDS is to solve the equations of motion Eq. (2), which can face a numerical complication due to the singularity of the matrix $M$ as defined in Eq. (4); that is, there exist non-zero vectors $x$ that satisfy $Mx=0$. In Appendix A, we show that Eq. (2) remains solvable since the vector $V$ given in Eq. (5) is always orthogonal to the null space of $M$. In this scenario, Eq. (2) has infinitely many solutions that lie on a hyperplane in the $N_{\boldsymbol{\theta}}$-dimensional parameter space. We show in Fig. 2 (a) a schematic of the hyperplane in a 2D space, which reduces to a straight line. Obviously one cannot solve Eq. (2) by directly inverting $M$. Instead, one could try to minimize $\norm{M\dot{\boldsymbol{\theta}}-V}^{2}$ using least-square (LSQ) techniques Endo et al. (2020b). An unbounded LSQ search can in principle reach any solution on the hyperplane, which can have a large magnitude. Consequently, a tiny time increment $\Delta t$ must be taken so that the changes in the rotation angles $\Delta\theta_{\mu}=\dot{\theta}_{\mu}\Delta t$ remain small enough to guarantee a continuous integration of the equations of motion. A remedy for this problem is to set a bound $b$ for the LSQ method, so the search is restricted to be within a hypercube of side $2b$ centered at the origin. An exact solution can still be found as long as the hypercube intersects with the hyperplane [see Fig. 2 (a)].
Tikhonov regularization Hanke and Hansen (1993) is another popular method for solving linear systems of equations of the form (2) with singular or ill-conditioned $M$. In this method, the solution to Eq. (2) is computed as $\dot{\boldsymbol{\theta}}=(M+\varepsilon I)^{-1}V$, where $\varepsilon$ is a small number and $I$ is the identity matrix. To better understand the effect of Tikhonov regularization, we first diagonalize the symmetric matrix $M$ with a unitary transformation $M=U\Lambda U^{\dagger}$, where $\Lambda$ is a diagonal matrix containing the eigenvalues $\lambda_{\mu}$ of $M$ in ascending order and where each column of $U$ gives the corresponding eigenvector. The eigenvalues $\lambda_{\mu}\geq 0$ since $M$ is positive semi-definite (see Appendix A). Assume the null space of $M$ has dimension $N_{\text{null}}$; then, $\lambda_{\mu}=0$ for $1\leq\mu\leq N_{\text{null}}$, and the null space of $M$ is spanned by the first $N_{\text{null}}$ eigenvectors of $M$: $\text{Null}(M)=\text{Span}\{u_{1},u_{2},\cdots,u_{N_{\text{null}}}\}$ where $u_{\mu}$ is the $\mu^{\text{th}}$ column of $U$. Under Tikhonov regularization, $\dot{\boldsymbol{\theta}}=(M+\varepsilon I)^{-1}V=U(\Lambda+\varepsilon I)^{-1}U^{\dagger}V$. $(\Lambda+\varepsilon I)^{-1}$ is also a diagonal matrix dominated by the first $N_{\text{null}}$ diagonal elements ($\varepsilon^{-1}$). However, since $V\perp\text{Null}(M)$, the first $N_{\text{null}}$ elements of $U^{\dagger}V$ are zero. Thus, as long as $\varepsilon$ remains large compared to the machine accuracy, the first $N_{\text{null}}$ elements of the resulting vector $(\Lambda+\varepsilon I)^{-1}U^{\dagger}V$, which gives the projection of $\dot{\boldsymbol{\theta}}$ onto $\text{Null}(M)$, are essentially zero. Meanwhile, the Tikhonov parameter $\varepsilon$ should also be much smaller than the nonzero eigenvalues $\lambda_{\mu}>0$. Thus, $\dot{\boldsymbol{\theta}}$ calculated in this way is close to the unique solution of Eq. (2) that is orthogonal to $\text{Null}(M)$. Such a solution is also schematically shown in Fig. 2 (a).
Alternatively, the coefficient matrix $M$ can be regularized by performing a truncation on its eigenvalues Hansen (1990). After diagonalizing $M$, Eq. (2) can be rewritten as $\Lambda(U^{\dagger}\dot{\boldsymbol{\theta}})=U^{\dagger}V$. Since the first $N_{\text{null}}$ elements of both $\lambda_{\mu}$ and $U^{\dagger}V$ are zero, a solution of the above linear system of equations can be straightforwardly obtained by letting $(U^{\dagger}\dot{\boldsymbol{\theta}})_{\mu}=0$ for $1\leq\mu\leq N_{\text{null}}$, and $(U^{\dagger}\dot{\boldsymbol{\theta}})_{\mu}=(U^{\dagger}V)_{\mu}/\lambda_{\mu}$ for $N_{\text{null}}<\mu\leq N_{\boldsymbol{\theta}}$. In other words, $U^{\dagger}\dot{\boldsymbol{\theta}}$ is an $N_{\boldsymbol{\theta}}$-dimensional vector whose first $N_{\text{null}}$ elements are zeros. After $U^{\dagger}\dot{\boldsymbol{\theta}}$ is obtained, $\dot{\boldsymbol{\theta}}$ is readily available as $\dot{\boldsymbol{\theta}}=U(U^{\dagger}\dot{\boldsymbol{\theta}})$. This procedure again leads to the unique solution located in the subspace orthogonal to $\text{Null}(M)$ as shown in Fig. 2 (a). In practice, $N_{\text{null}}$ is identified by truncating the eigenvalues based on a threshold value $\varepsilon$; in other words, $N_{\text{null}}=\left|\{\lambda_{\mu}|\lambda_{\mu}\leq\varepsilon\}\right|$ where ||$\cdot$|| ⋅ | denotes the size of a set. On noiseless simulators, this method can be interpreted as a modified Tikhonov regularization, where the uniform shift of the eigenvalues of $M$ by $\varepsilon I$ is replaced with a partial shift involving only the negligible eigenvalues within the null space. However, when noise is present, the first $N_{\text{null}}$ eigenvalues of $M=M_{0}+\delta M$ will take nonzero values $\delta\lambda_{\mu}$. Here, $M_{0}$ and $\delta M$ denote the exact coefficient matrix and the noise, respectively. At the same time, $u_{\mu}^{\dagger}V$ ($\mu\leq N_{\text{null}}$) will not vanish either, and their values $\delta(u_{\mu}^{\dagger}V)$ are also determined by the noise. Consequently, the projection of $\dot{\boldsymbol{\theta}}$ onto $\text{Null}(M_{0})$, calculated within Tikhonov regularization via $(\delta\lambda_{i}+\varepsilon)^{-1}\delta(u_{i}^{\dagger}V)$, can be large because it involves division of small quantities. On the other hand, such operations are largely avoided in the truncation method by effectively truncating the eigenvalues $\delta\lambda_{\mu}$ that are only nonzero due to noise, resulting in a solution in the vicinity of the orthogonal subspace of $\text{Null}(M_{0})$. This method, which we call the “truncation” method, has been applied to variational quantum imaginary time evolution Gacon et al. (2023). Here, we examine its performance on simulating real-time quantum dynamics on both noiseless and noisy simulators.
Figure 2: (a) Schematic showing different approaches for solving Eq. (2) in a 2D parameter space. The red dashed line shows the null space of $M$ (the collection of all $x$ satisfying $Mx=0$). The solid black line parallel to $\text{Null}(M)$ is the hyperplane containing all the solutions of Eq. (2). The black circles show the solutions that can be obtained using different methods. In particular, the Tikhonov regularization and the truncation methods can locate the unique solution orthogonal to $\text{Null}(M)$. (b) The growth of the number of unitaries $N_{\boldsymbol{\theta}}$ as a function of time, (c) the time evolution of the wavefunction infidelity, and (d) the time increment $\Delta t=\Delta\theta_{m}/\max_{\mu}\absolutevalue{\theta_{\mu}}$ | at each time step as a function of time, during the quench dynamics simulation of the TFIM with $N_{q}=12$, using the methods outlined in (a) to solve Eq. (2).
We compare the above-mentioned methods for solving the equations of motion Eq. (2) by performing AVQDS(T) on the TFIM with $N_{q}=12$. We here use method 3 of Fig. 1 to construct the compressed circuits. $\varepsilon$ is set to $10^{-6}$ for both the Tikhonov regularization and truncation methods. For the bounded LSQ search, we set the bound $b=5$. The variational parameters $\boldsymbol{\theta}$ are updated according to the Euler method $\Delta\boldsymbol{\theta}=\dot{\boldsymbol{\theta}}\Delta t$ with $\Delta t=\Delta\theta_{m}/\max_{\mu}\absolutevalue{\theta_{\mu}}$ | Iserles (2009), where the maximal allowed change in rotation angles $\Delta\theta_{m}$ is set to 0.005. Fig. 2 (b) plots the growth of the number of unitaries $N_{\boldsymbol{\theta}}$ as a function of $t$, which shows that the different methods produce circuits of similar complexity. We show the infidelity $1-f$ of |$\ket{\Psi[\boldsymbol{\theta}(t)]}$ and the time increment at each time step ($\Delta t$) as a function of $t$ in Fig. 2 (c) and (d), respectively. The fidelity $f$ is defined as $f\equiv\absolutevalue{\langle\Psi[\boldsymbol{\theta}(t)]|\,\Psi_{\text{exact}}(t)\rangle}^{2}$, where |$\ket{\Psi_{\text{exact}}(t)}=e^{-i\hat{\mathcal{H}}t}\ket{\Psi_{0}}$ is obtained by directly applying the time-evolution operator. All the curves coincide for $t<2$ in both Fig. 2 (c) and (d), indicating the matrix $M$ is well-conditioned, and all the methods can effectively locate the unique solution to Eq. (2) during this time period.
In the current NISQ era, it is important to examine the noise resilience of different strategies for solving Eq. (2) in a noisy environment. Here, we only consider the shot noise on the matrix $M$—resulting from a fixed set of circuit samples or “shots” performed per matrix element—to single out the effect of these solvers on regularizing $M$. The matrix elements of $M$ can be obtained by calculating the probability $p$ that an ancilla qubit measurement yields 1 Yao et al. (2021), such that $M_{\mu\nu}=(2p-1)/4$. The variance $\sigma_{M_{\mu\nu}}^{2}=p(p-1)/(4N_{s})$ where $N_{s}$ is the number of shots used to measure $M_{\mu\nu}$. We assume that all matrix elements are measured with the same number of shots. In our calculations, we simulate the shot noise by replacing the noiseless value $M_{\mu\nu}$ with a random number drawn from a Gaussian distribution with mean $M_{\mu\nu}$ and variance $\sigma_{M_{\mu\nu}}^{2}$.
Since the bounded LSQ method requires a predetermined bound that is generally problem-specific, we focus on the Tikhonov regularization and truncation methods, taking the TFIM with $N_{q}=10$ as an example. A fixed HVA is adopted to ensure that the quantum circuits are the same throughout the simulation for both methods. Relatively deeper circuits are required in the presence of noise; thus, we choose $L=15$ and 20. $N_{s}$ is set to moderate values $10^{4}$ and $10^{5}$. The parameter $\varepsilon$ controls the small positive value added to the diagonal of the matrix $M$ in the Tikhonov regularization method, or the threshold for truncating the eigenvalues of $M$ in the truncation method. Relative to noiseless simulations, a larger $\varepsilon$ is needed so as to avoid being overwhelmed by the noise. On the other hand, if $\varepsilon$ is too large, it causes too much distortion to the matrix $M$. After some pre-experimentation, we choose $\varepsilon=10^{-2}$ and $10^{-3}$ for the Tikhonov regularization method, and $\varepsilon=10^{-3}$ and $10^{-4}$ for the truncation method. These parameters yield the same order of magnitude for the standard error of $M_{\mu\nu}$ given $N_{s}=10^{4}$ and $10^{5}$. We implement a fixed time step $\Delta t$ for this comparison so that the total computation cost is the same for both methods. $\Delta t$ is set to 0.002 for $L=15$ and 0.005 for deeper circuits with $L=20$. 20 independent runs were performed for each parameter set. Figure 3 gives the time traces of the wavefunction infidelity, in which the region shaded in light purple shows the “acceptance range” with $f>0.9$. The contrast is clear: the simulations using Tikhonov regularization escape the acceptance range shortly after being launched, while those performed with the truncation method stay within the acceptance range for a much longer time, which increases with the number of HVA layers or the number of shots. The failure of the Tikhonov regularization method is due to its difficulty in curbing the components of $\dot{\boldsymbol{\theta}}$ on the null space of the unperturbed coefficient matrix, resulting in large $|\dot{\boldsymbol{\theta}}|$. One has to reduce $\Delta t$ by about an order of magnitude in order to make it work, greatly increasing the required quantum resources. The truncation method with $\varepsilon=10^{-3}$ works the best for all the parameter sets that have been examined, and will be used in subsequent noisy simulations.
IV.2 Benchmarking AVQDS(T) on noiseless simulators
@@
In Fig. 5, we plot the evolution of the wavefunction infidelity $1-f$, the circuit depth, and the number of CNOT gates during quench dynamics simulations of the three models. The AVQDS and AVQDS(T) methods generate similar wavefunction infidelities, both smaller than 1%, for all three models (see top panel of Fig. 5). On the other hand, for all three systems the circuit depth at the end of the AVQDS(T) simulations is reduced by more than half relative to AVQDS (see middle panel of Fig. 5). This reduction is important as the circuit depth determines the total execution time of the circuit on hardware, which is upper bounded by the coherence time of the device. Compressing the circuits thus allows reaching a larger final simulation time on hardware. While the total number of unitaries in the final circuit is similar for AVQDS(T) and AVQDS, the number of resource-intensive two-qubit gates is reduced by 22% and 26% for the TFIM ($N_{q}=14$) and the MFIM ($N_{q}=12$), respectively, resulting in a noticeable reduction in the number of CNOT gates, as can be seen in Fig. 5 (c) and (f). Such an effect is inapplicable to the HM ($N_{q}=10$) since we choose to only include two-qubit operators in HM simulations. In contrast, if the standard Trotter decomposition technique is used to perform the dynamical simulations, the circuit depth and the number of CNOTs must increase by over an order of magnitude compared with AVQDS(T) in order to reach similar accuracy. HVA exemplifies a group of methods with a fixed quantum circuit depth. As shown in the top panel of Fig. 5, HVA can deliver the required accuracy in the early stages of the simulations. However, after a certain period of time, the applied HVA circuits completely fail to describe the quantum dynamics as the wavefunction infidelity increases rapidly and approaches one. Deeper circuits would have been required at this late stage. This behavior, observed in all the models we have studied, demonstrates a typical disadvantage of using a fixed-depth ansatz for quantum-dynamical simulations: one either uses a deep ansatz for the entire duration of the simulation (even when it is not needed to represent the time evolved state at early times) or loses accuracy at long times. On the other hand, with the ability to expand the ansatz on demand, AVQDS(T) represents a much more efficient strategy for allocating quantum resources.

### v2

@@ III Models
In addition to AVQDS and AVQDS(T), we also simulated quantum dynamics with the first-order Trotter decomposition [50, 48] and VQDS with a fixed ansatz, namely the Hamiltonian variational ansatz (HVA) [51]. In Trotter decomposition, the time evolution of a quantum state over a single time step $\delta t$ is approximated as |$\ket{\Psi(t+\delta t)}\approx\prod\limits_{\mu}e^{-i\delta tj_{\mu}\,\hat{h}_{\mu}}\ket{\Psi(t)}$, where $j_{\mu}\hat{h}_{\mu}$ are the terms in the Hamiltonian: ^$\hat{\mathcal{H}}=\sum\limits_{\mu}j_{\mu}\hat{h}_{\mu}$. $\hat{h}_{\mu}$ are usually Pauli strings and $j_{\mu}$ are constants. A basic feature of such Trotter circuits is that the circuit depth keeps increasing with the simulation time $t$ as the circuit depth scales linearly with $t$. Motivated by the form of these Trotter circuits, the HVA entails parameterized unitary gates generated by the individual Pauli terms $\hat{h}_{\mu}$. The ansatz uses a fixed number of layers, $L$, and the rotation angles entering the parameterized unitaries are allowed to vary in space and from layer to layer. Thus, the variational quantum state within HVA can be written as
@@ IV Results and Discussions
IV.1 Benchmarking AVQDS(T) on noiseless simulators
@@
In Fig. 3, we plot the evolution of the wavefunction infidelity $1-f$, the circuit depth, and the number of CNOT gates during quench dynamics simulations of the three models. The AVQDS and AVQDS(T) methods generate similar wavefunction infidelities, both smaller than 1%, for all three models (see top panel of Fig. 3). On the other hand, for all three systems the circuit depth at the end of the AVQDS(T) simulations is reduced by more than half relative to AVQDS (see middle panel of Fig. 3). This reduction is important as the circuit depth determines the total execution time of the circuit on hardware, which is upper bounded by the coherence time of the device. Compressing the circuits thus allows reaching a larger final simulation time on hardware. While the total number of unitaries in the final circuit is similar for AVQDS(T) and AVQDS, the number of resource-intensive two-qubit gates is reduced by 22% and 26% for the TFIM ($N_{q}=14$) and the MFIM ($N_{q}=12$), respectively, resulting in a noticeable reduction in the number of CNOT gates, as can be seen in Fig. 3 (c) and (f). However, we believe this reduction is problem-dependent, and can hardly be generalized to an arbitrary Hamiltonian. Such an effect is inapplicable to the HM ($N_{q}=10$) since we choose to only include two-qubit operators in HM simulations.
In contrast, if the standard Trotter decomposition technique is used to perform the dynamical simulations, the circuit depth and the number of CNOTs must increase by over an order of magnitude compared with AVQDS(T) in order to reach similar accuracy. HVA exemplifies a group of methods with a fixed quantum circuit depth. As shown in the top panel of Fig. 3, HVA can deliver the required accuracy in the early stages of the simulations. However, after a certain period of time, the applied HVA circuits completely fail to describe the quantum dynamics as the wavefunction infidelity increases rapidly and approaches one. Deeper circuits would have been required at this late stage. This behavior, observed in all the models we have studied, demonstrates a typical disadvantage of using a fixed-depth ansatz for quantum-dynamical simulations: one either uses a deep ansatz for the entire duration of the simulation (even when it is not needed to represent the time evolved state at early times) or loses accuracy at long times. On the other hand, with the ability to expand the ansatz on demand, AVQDS(T) represents a much more efficient strategy for allocating quantum resources.
IV.2 Implementation of realistic noise channels
In practical quantum computing, NISQ hardware is subject to various error sources including coherent errors caused by imperfect gate operations and stochastic errors due to qubit decoherence, dephasing, and relaxation. Here we investigate how these hardware imperfections affect quantum dynamical simulations using the TFIM with $N_{q}=6$ as an example. The final converged ansatz obtained in a noiseless simulation is used in these noisy simulations. We implement a phenomenological noise model proposed by Kandala et al[19], which consists of an amplitude damping channel ($\Lambda_{a}[\rho]=\sum_{i=1}^{2}E_{i}^{a}\rho E_{i}^{a\dagger}$) and a dephasing channel ($\Lambda_{d}[\rho]=\sum_{i=1}^{2}E_{i}^{d}\rho E_{i}^{d\dagger}$). Here, $\rho$ is the qubit density matrix. The Kraus operators are defined as follows:
$E_{1}^{a}$ $=$ $\begin{pmatrix}1&0\\ 0&\sqrt{1-p^{a}}\end{pmatrix},E_{2}^{a}=\begin{pmatrix}0&\sqrt{p^{a}}\\ 0&0\end{pmatrix},$
$E_{1}^{d}$ $=$ $\begin{pmatrix}1&0\\ 0&\sqrt{1-p^{d}}\end{pmatrix},E_{2}^{d}=\begin{pmatrix}0&0\\ 0&\sqrt{p^{d}}\end{pmatrix}.$ (10)
The error rates $p^{a}=1-e^{-t_{g}/T_{1}}$ and $p^{d}=1-e^{-2t_{g}/T_{\phi}}$ depend on the gate time $t_{g}$, the qubit relaxation time $T_{1}$, and the dephasing time $T_{\phi}=2T_{1}T_{2}/(2T_{1}-T_{2})$, where $T_{2}$ is the qubit coherence time. For simplicity, we use a uniform single-qubit gate error rate $p_{1}^{a}=p_{1}^{d}\equiv p_{1}=10^{-4}$, which is typical of the current hardware. The two-qubit gate error is also assumed to be uniform $p_{2}^{a}=p_{2}^{d}\equiv p_{2}$, with $10^{-4}<p_{2}<10^{-2}$ [52, 53]. We used $2^{14}$ shots for all measurements. The calculations were performed using the QASM simulator provided in the quantum computing package Qiskit [54].
[…]
In Fig. 4, we show the results of the noisy simulations with the two-qubit error rate $p_{2}=10^{-4},10^{-3}$, and $10^{-2}$, together with the results obtained on a noiseless simulator. The wave function fidelity progressively worsens as $p_{2}$ increases compared with the noiseless results, as can be seen in Fig. 4 (a). The light purple region in Fig. 4 (a) shows an “acceptance range” with $f>0.9$. For $p_{2}=10^{-4}$ and $10^{-3}$, the fidelity is largely within this range in the entire duration of $t<2$, while the simulation with $p_{2}=10^{-2}$ leaves the acceptance range at $t\sim 0.5$. The measured total energy is plotted in Fig. 4 (b). For this closed system, the exact total energy is kept constant at $E=-6$, which is accurately captured in the noiseless simulation. For $p_{2}=10^{-4}$ and $10^{-3}$, the measured total energy gradually deviates from the exact value during the simulation, with the error growing to 7% and 12% at $t=2$, respectively. For $p_{2}=10^{-2}$, the total energy oscillates around the exact value with the largest deviation exceeding 20%. We also measured the $X$-component of the magnetization defined as $S_{X}=\bra{\Psi(t)}\sum_{i=1}^{N_{q}}X_{i}\ket{\Psi(t)}$, as shown in Fig. 4 (c). For both $p_{2}=10^{-4}$ and $p_{2}=10^{-3}$, the trajectory of $S_{X}$ generally follows that of the noiseless simulation, with $p_{2}=10^{-4}$ giving an overall better agreement. For $p_{2}=10^{-2}$, on the other hand, $S_{X}$ starts to develop a large deviation from the noiseless values at $t=0.6$.
IV.3 Comparison with the adaptive pVQD method
The projected variational quantum dynamics (pVQD) method [37] and its adaptive version [38] represent an alternative approach for simulating quantum dynamics using adaptively parameterized quantum circuits. In pVQD, the variational parameters $\boldsymbol{\theta}$ are updated at each time step by maximizing the overlap between the parametrized state at the next time step |$\ket{\Psi[\boldsymbol{\theta}+\boldsymbol{\delta\theta}]}$ and the Trotter-evolved state from the current parameters $\hat{\mathcal{T}}(\delta t)\ket{\Psi[\boldsymbol{\theta}]}$, where the Trotter operator $\hat{\mathcal{T}}$ is an approximation of the exact time evolution operator $\hat{\mathcal{T}}(\delta t)\approx e^{-i\hat{H}\delta t}$, and $\delta t$ is a predetermined time-step size. As pointed out in Ref. [37], in the limit of $\delta t\rightarrow 0$, pVQD is equivalent to the McLachlan variational principle implemented in AVQDS(T).
It will bring useful insight into the general performance of variational quantum algorithms in simulating quantum dynamics by comparing these two methods. This comparison naturally incorporates the impact of time-step size. For simplicity, we assume a constant number of shots for each distinct circuit used to measure quantities of interest. For example, the same number of shots is applied to measure a quantum state overlap in adaptive pVQD and an element of the quantum geometric tensor in AVQDS. This assumption allows us to focus solely on the number of distinct circuits for comparison.
In AVQDS [34], the number of distinct circuits for measuring $M$, $V$ and $L^{2}$ at each time step can be estimated as $N_{H}(N_{\boldsymbol{\theta}}-1)+N_{\boldsymbol{\theta}}(N_{\boldsymbol{\theta}}+1)/2+N_{H,H^{2}}^{c}$, where $N_{\boldsymbol{\theta}}$ is the number of variational parameters, $N_{H}$ is the number of Pauli strings in the Hamiltonian $\hat{H}$, which scales linearly with $N_{q}$ for the local spin models studied in this paper. The constant $N_{H,H^{2}}^{c}$ is the number of partitions of the commuting Pauli strings in $\hat{H}$ and $\hat{H}^{2}$. If the ansatz adaptive process is triggered, it requires additional $N_{H}N_{\boldsymbol{\theta}}$ circuits for scanning the operator pool composed of the $N_{H}$ non-identity Pauli strings in the Hamiltonian. For the adaptive pVQD method, the dominant computational cost is the measurement of gradients using the parameter shift rule during the scan of the pool and the optimization of the cost function. Accordingly, the number of distinct circuits at each time step can be estimated as $2N_{\boldsymbol{\theta}}^{0}N_{\mathrm{iter}}^{0}+\sum_{i=1}^{N_{\mathrm{adapt}}}(2N_{H}+2N_{\boldsymbol{\theta}}^{i}N_{\mathrm{iter}}^{i})$, where $N_{\mathrm{adapt}}$ is the total number of adaptive steps in the current time step (which can be zero), $N_{\boldsymbol{\theta}}^{i}$ and $N_{\mathrm{iter}}^{i}$ are the number of parameters and the number of iterations for optimizing the cost function after the $i$th adaptive step, respectively.
To illustrate this comparison concretely, we analyze quench dynamics simulations of the transverse-field Ising model (TFIM) with $N_{q}=10$ sites. The adaptive pVQD calculations were performed using the open-source code [55] with the default settings for trotterization (second order with 2 repetitions), which allows for relatively large time-step sizes. Specifically, $\delta t=0.16$ gives the most accurate result for simulations run up to $t=2$, with a minimal final infidelity $0.0121-f=$ 0.012. $7.78\times 10^{5}$ distinct circuits are required for quantum measurements during the simulation. For AVQDS, a much smaller time-step size of 0.02 is necessary to achieve similar infidelity of $0.0101-f=$ 0.010 at $t=2$. However, the cumulative total number of distinct circuits is $7.31\times 10^{5}$, which is comparable to the adaptive pVQD method with $\delta t=0.16$. The final circuit depth for the variational ansatz in adaptive pVQD and AVQDS simulations is 22 for both methods. For AVQDS, the most complicated circuit is represented by measuring the element of the quantum geometric tensor involving the final ansatz parameter, which introduces an ancillary qubit and enlarges the circuit depth to $22+4=26$, due to the controlled two-qubit Pauli gates. While in adaptive pVQD, one needs to double the ansatz circuit plus an extra trotter circuit (second order with two repetitions) for measuring the cost function, which makes the circuit depth $22\times 2+(4\times 3-3)=53$ (each Trotter layer has a depth of 3 for the TFIM).